OpenStudy (anonymous):

Sally was asked to find two integers such that one is three less than 4 times the other and the sum of the two integers is 42. The integers must be a 14,28 b 15, 27 c 9,33 d -13,55

OpenStudy (anonymous):

Lets call the integers a & b a is three less than 4* b: a + 3 = 4b and a + b = 42 Can you solve the system of these two equations using substitution?

OpenStudy (anonymous):

would i subtract 42?

OpenStudy (anonymous):

im just really confused

OpenStudy (anonymous):

a + b = 42 ==> a = 42 - b then you can substitute: a + 3 = 4b ==> 42 - b + 3 = 4b Can you simplify and solve this for b? Once you have b, plug it into a + b = 42 to find a.